OpenStudy (anonymous):

Let A be a square matrix with A^2 = 0. Show that I + A is invertible and (I + A)^−1 = I − A. Use this result to find B^−1, where B = (1 2 3 0 1 0 0 0 1)

OpenStudy (turingtest):

\[(I-A)(I+A)=?\]

OpenStudy (anonymous):

I don't know what the dimensions of A would be. If I knew what A is I could find its inverse.

OpenStudy (turingtest):

you do not need the dimensions of A just do the multiplication and what do you get?

OpenStudy (anonymous):

I^2 - A^2?

OpenStudy (turingtest):

yeah, and what are they telling us that A^2 equals?

OpenStudy (anonymous):

0...

OpenStudy (turingtest):

so we can write\[(I-A)(I+A)=I\]this implies that for the matrix \(B=I+A\) there exists another matrix \(B^{-1}\) such that\[B^{-1}B=I\]which is the definition of an invertible matrix hence \(B=I+A\) is invertible, and it's inverse is \[B^{-1}=(I+A)^{-1}=I-A\]

OpenStudy (turingtest):

is this making any sense at all to you?

OpenStudy (anonymous):

Yes. I know that B^-1(B) will give you an identity matrix.. Now I have to solve for B^-1.

OpenStudy (turingtest):

well, notice that \(B^2\neq0\), but if we split B up we have \(B=A+I\) where \(A^2=0\) now we can just apply the theorem we just proved